# Revisiting MAE pre-training for 3D medical image segmentation

### v2

@@ 1 Introduction
P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images [44, 42, 30, 31, 45, 12], almost approaching scales of supervised dataset sizes. These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire. While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns. Some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Decline (ABCD) dataset of the NIH (N=40k) or the UK Biobank (UK-BB) (N=120k), but they restrict access pending an internal review board’s approval, posing a hurdle for open science.
P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers [42, 31, 39, 5]. While transformers are prevalent in the 2D natural imaging domain [8], recent architectures leveraging attention [36] have so far not been able to reach state-of-the-art performance in 3D medical segmentation. In fact, well-configured 3D U-Net [6, 27] inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin [19]. This underscores the need for SSL methods that can be seamlessly integrated with CNNs to harness their full potential in medical image analysis on downstream tasks.
P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy. This is represented through: i) Evaluating on too few datasets to show generalization of the pre-training ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture [39] iii) Comparing to bad from-scratch baselines, like a badly configured, outdated [6] model instead of a well-configured 3D nnU-Net CNN baseline [18]. iv) Evaluating their method on seen data they pre-trained on. We want to emphasize that we do not intend to point fingers, but to raise awareness that evaluation matters and insufficient evaluation can lead to a lack of clarity which methods are the best. This observation is similar to the recent study by Isensee et al. [19], that shows that this is also prevalent in the field of medical image segmentation when training from-scratch.
@@
    State-of-the-Art Performance: Our final model achieves superior performance compared to existing SSL-pre-trained models using the same architecture and dataset, see Fig. 1. Notably, this is the first work to demonstrate that SSL pretraining with a fixed architecture can consistently outperform a state-of-the-art, dynamically optimized nnU-Net baseline by approximately 3 Dice points across 11 diverse downstream datasets [18]. An associated model will be made available.
@@
    Comprehensive Evaluation Addressing Previous Pitfalls: Building on the detailed analysis of limitations in prior work (see Tab. 1), we set up a robust development environment to design our method. () P1: We utilize a collection of 39k 3D MRI volumes for self-supervised pre-training, substantially larger than most datasets used in prior SSL work [44, 42, 30, 31, 45, 12]. () P2: By using the state-of-the-art Residual Encoder U-Net CNN architecture [19], we bypass the limitations of outdated or transformer-based architectures, leveraging CNNs’ superior performance in 3D medical segmentation. () P3: We rigorously evaluate our model across five development and eight testing datasets, covering a wide range of downstream tasks, diverse pathologies, novel image modalities, and multi-center datasets. Moreover, we disentangle pre-training methodology from other contributions like novel backbones and include the strong 3D nnU-Net baseline, ensuring reliable results.
@@ 2 Development Framework
The goal of this paper is to develop a robust SSL pretraining method. Due to the limited prior work in the 3D medical domain, many design choices need to be made. We address this by sequentially validating each methodological contribution on five downstream development datasets before testing the final configuration on eight untouched test datasets. To reduce the search space and disentangle the effects of SSL pre-training from other basic design choices, we choose to keep some parameters fixed based on best practices in the domain:
@@
Additionally, eight hold-out test sets were used to evaluate the efficacy of our learned representations when fine-tuning them to segment other target structures.
@@
For test datasets D6-D12 we use an 80/20 split for fine-tuning and for testing. A validation set is omitted as we fine-tune each method only once without any interventions. For D13 all data is used for testing, as no training is conducted.
@@ 4 Revisiting 3D MAEs
(a) Sparsification: The introduction of all sparsification components showed best results, with most improvements from adding the additional densification conv layer between encoder and decoder.
@@
(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios. Overall, a dynamic range, including higher masking ratios, performs the best.
@@
SSL Method No (Dyn.) No (Fix.) S3D* S3D
Dataset Dice Similarity Coefficient (DSC)
MS FLAIR (D1) 57.81 59.82 59.75 60.35
Brain Mets (D2) 63.66 56.53 64.20 65.24
Hippocampus (D3) 89.18 89.24 89.45 89.60
Atlas22 (D4) 63.28 65.52 66.61 66.95
CrossModa (D5) 85.64 83.44 83.61 84.08
Cosmos22 (D6) 60.28 78.17 80.01 80.00
ISLES22 (D7) 77.94 79.44 78.94 79.70
Hanseg (D8) 59.00 61.85 61.27 62.11
HNTS-MRG24 (D9) 66.73 65.90 67.03 68.62
BRATS24 Africa (D10) 93.07 92.51 92.49 92.19
Avg. DSC 71.66 73.24 74.34 74.88
Avg. Rank 3.2 2.9 2.4 1.5
@@
The masked region is determined by sampling randomly in the CNN’s bottleneck of shape [5x5x5] and up-sampling these regions to the input resolution, to ensure the masks align in the bottleneck of the CNN architectures. This results in masking regions of [32x32x32] voxels of non-overlapping regions in the input. As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos [14, 9]. In the scope of the development phase we explore 5 static masking ratios between 30 and 90%, and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
Results are presented in Table Tab. 2(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well. Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn and with the upcoming scaling experiment in mind. We refer to this model as Spark3D (S3D) going forward.
@@
Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning? We investigated various different schedules, with all of themimplemented in the nnU-Net framework [18]. Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder. Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR. When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder. In some configurations, this results in two learning rate warm-ups of 12.5k steps each. Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well. Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
Results are presented in Table 3 and allow three important observations to be made: (i) Warm-up stages are essential:Not applying a warm-up step significantly reduces performance. Including a warm-up for both the encoder and decoder boosts accuracy by 0.6 to 1 DSC points. (ii) Learning rate adjustments matter:Reducing the peak learning rate to 1e-3 during fine-tuning consistently yields better results than the default 1e-2, with the best performance seen when fine-tuning both the encoder and decoder with lower learning rates. (iii) Freezing encoder weights is detrimental:The encoder should not remain fixed during fine-tuning. Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned.
@@ 5 Results and Discussion
We compare our final model S3D against Volume Contrastive (VoCo) [42], VolumeFusion (VF) by Wang et al. [39], as well as Models Genesis (MG) [44]. The baselines are pre-trained for 250k steps using the same framework on the same data with the same backbone and the same hyperparameters (where possible) but scaled to fully utilize an A100 40GB GPU. We provide explicit baseline methods and configuration details in Appendix A. Moreover, we compare against two strong from-scratch baselines. The first, No (Dyn.), represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings. The second from-scratch baseline, No (Fixed), is a nnU-Net training with the same plans and pre-processing as defined by our pre-training. The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Tab. 4. Additionally, the ranking stability of the methods is evaluated through bootstrapping and is provided in Fig. 2.
@@ 5.1 Observations
Across all tested datasets, SSL pre-trained methods demonstrate improved downstream segmentation performance. Comparing our S3D method to the most similar from-scratch baseline No (Fixed), we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points. This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and state-of-the-art architecture.
MAEs dominate
Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG and S3D) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training. Given the age of Models Genesis - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF. We attribute this to a combination of two factors: 1. Models Genesis was originally published and trained on an outdated 3D-UNet [6] and outside of the powerful nnU-Net framework [18]. This highlights the importance of avoiding Pitfall 2: Training on a state-of-the-art backbone. 2. VoCo and VF were introduced in conjunction with architectures they were optimized for. By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture pre-training combination may be suboptimal for the new CNN backbone.
@@ 5.2 Ablation experiments
To assess the generalization capability of the proposed pre-training method, we tested two scenarios. First, we evaluated fine-tuning our method on an unseen modality using the TOF Angiography Aneurysms dataset (D12). As shown in Tab. 9 without pre-training, the fixed configuration suffers a performance drop of 20 Dice points. We attribute this to the significant difference in median spacing for the downstream task ($[0.50,0.43,0.43]$ mm), which has a higher resolution than the fixed target spacing of $[1,1,1]$ mm used in the pre-training experiments. This lower resolution likely increases the difficulty of segmenting small aneurysms. Despite this decrease, pre-training mitigates some of this degradation and proves highly beneficial compared to training the same configuration from scratch. Interestingly, ModelGenesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences. Second, we fine-tuned on the D2 dataset using only the T1 contrast-enhanced (T1ce) sequence and applied these models directly to D13 without any additional fine-tuning. While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that MAE pre-training improves generalization across different centers, with our S3D model yielding the best performance.
@@ Appendix B Longer Training schedule
Table 5: Public available checkpoint trained on the ABCD dataset: To provide a public available checkpoint, we retrained our proposed model on the ABCD dataset, indicated by a *. It performs slightly worse than the network pre-trained on the private dataset.
@@ C.2 Comparison to previous work on CT data
Table 6: Forty images with SSL are almost as good as all data from-scratch! S3D model almost reaches the performance of the model trained from-scratch with only 40 training cases. * D1 has only 38 training cases for the train split.
@@
Experiment Setting No Dyn. No Fixed VoCo VF MG S3D-B S3D-L
Modality shift TOF Angio. Aneurysms(D12) 42.61 22.76 22.32 31.21 34.60 28.72 26.90
In Distribution Brain Mets (D2) 72.81 67.93 64.34 71.69 69.05 71.56 72.53
Patient shift Brain Mets (D13) 64.08 61.61 56.78 63.95 64.22 64.54 64.95
Table 10: Fine-tuning length: When initializing from our pre-trained checkpoint, it is possible to achieve a large fraction of the final performance after less than 15% of the normal training time. Despite this a full training schedule reaches better performance.

### v3

@@ 1 Introduction
P1 - Limited dataset size: Many SSL approaches have been developed on very few unlabeled volumes, often being trained on fewer than 10,000 images [48, 46, 33, 34, 49, 14, 13, 3], almost approaching scales of supervised dataset sizes. These datasets tend to be pooled from publicly available annotated datasets, as larger datasets pose a greater hurdle to acquire. While many hospitals possess 3D medical images in the millions, they are locked away from the public due to patient privacy concerns. Some larger open-source datasets exist, e.g., the Adolescent Brain Cognitive Development (ABCD) dataset of the NIH (N=40k) [27] or the UK Biobank dataset (N=120k), but they restrict access pending an internal review board’s approval, posing a hurdle for open science.
P2 - Outdated Backbones: Many studies develop their SSL method on non-state-of-the-art architectures, e.g., utilizing transformers [46, 34, 43, 6, 13, 3]. While transformers are prevalent in the 2D natural imaging domain [9], recent architectures leveraging attention [40] have so far not been able to reach state-of-the-art performance in 3D medical segmentation. In fact, well-configured 3D U-Net [7, 30] inspired CNNs dominate 3D medical image segmentation, outperforming transformer-based models by a large margin [21]. This underscores the need for SSL methods that can be seamlessly integrated with CNNs to harness their full potential in medical image analysis on downstream tasks.
P3 - Insufficient Evaluation: Existing methods often lack rigorous evaluation, masking the methods’ efficacy. This is represented through: i) Evaluating too few datasets to show generalization of the pre-training [39, 3] ii) Stacking multiple contributions, e.g., novel architecture with new pre-training, which does not allow one to draw conclusions if the pre-training is effective without the architecture [43] iii) Comparing to bad from-scratch baselines, like a badly configured, outdated Çiçek et al. [7] UNet instead of a well-configured 3D nnU-Net CNN baseline [20]. iv) Evaluating their method on seen data they pre-trained on. We want to emphasize that we do not intend to point fingers, but to raise awareness that evaluation matters and insufficient evaluation can lead to a lack of clarity which methods are the best. This observation is similar to the recent study by Isensee et al. [21], which shows that this is also prevalent in the field of medical image segmentation when training from scratch.
@@
    State-of-the-Art Performance: Our final model achieves superior performance compared to existing SSL-pre-trained models using the same architecture and pre-training dataset, see Fig. 1. Notably, this is the first work to demonstrate that SSL pre-training with a fixed architecture can consistently outperform a state-of-the-art, dynamically optimized nnU-Net baseline by approximately 3 Dice points across 11 diverse downstream datasets [20].
@@
    Comprehensive Evaluation Addressing Previous Pitfalls: Building on the detailed analysis of limitations in prior work (see Tab. 1), we set up a robust development environment to design our method. () P1: We utilize a collection of 39k 3D MRI volumes for self-supervised pre-training, substantially larger than most datasets used in prior SSL work [48, 46, 33, 34, 49, 14, 39]. () P2: By using the state-of-the-art Residual Encoder U-Net CNN architecture [21], we bypass the limitations of outdated or transformer-based architectures, leveraging CNNs’ superior performance in 3D medical segmentation. () P3: We rigorously evaluate our model across five development and eight testing datasets, covering a wide range of downstream tasks, diverse pathologies, novel image modalities, and multi-center datasets. Moreover, we disentangle the pre-training methodology from other contributions like novel backbones and include the strong 3D nnU-Net baseline, ensuring reliable results.
@@ 2 Development Framework
The goal of this paper is to develop a robust SSL pre-training method. Due to the limited prior work in the 3D medical domain, many design choices need to be made. We address this by sequentially validating each methodological contribution on five downstream development datasets before testing the final configuration on eight untouched test datasets. To reduce the search space and disentangle the effects of SSL pre-training from other basic design choices, we choose to keep some parameters fixed based on best practices in the domain:
@@
Additionally, eight hold-out test datasets were used to evaluate the efficacy of our learned representations when fine-tuning them to segment other target structures.
@@
For test datasets D6-D12 we use an 80/20 split for fine-tuning and for testing. A validation set is omitted as we fine-tune each method only once without any interventions. For D13 all data is used for testing, as no training is conducted. D12 and D13 are only used to assess generalization performance, with D13 only appearing in Appendix Tab. 9.
@@ 4 Revisiting 3D MAEs
(a) Sparsification: The introduction of all sparsification components showed the best results, with most improvements coming from adding the additional densification conv layer between encoder and decoder.
@@
(b) Masking ratio: Masking ratios between 60% and 75% worked best when choosing static ratios. However, a dynamic range, including higher masking ratios, performs similarly and was chosen.
@@
Pre-training D1 D2 D3 D4 D5 Avg. D1-D5
No Dyn. 45.56 72.26 88.80 60.44 82.61 69.93
No Fixed 49.37 69.13 88.78 60.74 81.33 69.87
VoCo 50.35 67.20 88.22 57.82 80.29 68.77
VF 49.93 69.58 88.83 61.75 81.48 70.31
MG 50.50 71.14 88.83 63.29 82.15 71.18
S3D (ours) 51.49 74.01 88.83 62.39 81.54 71.65
@@
The masked region is determined by sampling randomly in the CNN’s bottleneck of shape [5x5x5] and up-sampling these regions to the input resolution, to ensure the masks align in the bottleneck of the CNN architectures. This results in masking regions of [32x32x32] voxels of non-overlapping regions in the input. As a sampling strategy, we follow random masking, as previous work showed no benefit of structured masking for images nor videos [16, 10]. In the scope of the development phase, we explore 5 static masking ratios between 30% and 90% and evaluate a dynamic masking ratio randomly masking between 60% and 90%.
Results are presented in Tab. 2(b) and highlight that the masking ratios of 60%, 75%, and the dynamic masking ratio of 60% to 90% perform equally well. Due to the highly similar performance, we choose to proceed with a dynamic masking ratio over the static masking ratio, due to expecting this masking to be more difficult to learn. We refer to this model as Spark3D (S3D) going forward.
@@
Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning? We investigated various different schedules, with all of them implemented in the nnU-Net framework [20]. Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder. Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR. When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder. In some configurations, this results in two learning rate warm-ups of 12.5k steps each. Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the warm-up process or to fine-tune the encoder weights as well. Lastly, we investigate whether to decrease the peak learning rate to 1e-3, 1e-4 or to keep it at the default of 1e-2.
Results are presented in Table 3 and allow three important observations to be made: (i) Warm-up stages are essential:Not applying a warm-up step significantly reduces performance. Including a warm-up for both the encoder and decoder boosts accuracy by 0.6 to 1 DSC points. (ii) Learning rate adjustments matter:Reducing the peak learning rate to 1e-3 during fine-tuning consistently yields better results than the default 1e-2, with the best performance seen when fine-tuning both the encoder and decoder with lower learning rates. (iii) Freezing encoder weights is detrimental:The encoder should not remain fixed during fine-tuning. Allowing the encoder to be fine-tuned improves the model’s performance compared to when only the decoder is fine-tuned, as observed in early exploratory work.
@@ 5 Results and Discussion
We compare our final model S3D against Volume Contrastive (VoCo) [46], VolumeFusion (VF) [43], as well as Models Genesis (MG) [48]. The baselines are pre-trained for 250k steps using the same framework on the same data with the same backbone and the same hyperparameters (where possible) but scaled to fully utilize an A100 40GB GPU. We provide explicit baseline methods and configuration details in Appendix A. Moreover, we compare against two strong from-scratch baselines. The first, No (Dyn.), represents a from scratch default nnU-Net 3D architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings. The second from-scratch baseline, No (Fixed), is a nnU-Net training with the same plans and pre-processing as defined by our pre-training. The dataset-wise mean dice similarity coefficient (DSC) and mean normalized surface distance (NSD) at 2 mm tolerance values across our test dataset suite are provided in Tab. 4. Additionally, the ranking stability is evaluated through bootstrapping and is provided in Fig. 2.
@@ 5.1 Observations
Across all tested datasets, SSL pre-trained methods demonstrate improved downstream segmentation performance. Comparing our S3D method to the most similar from-scratch baseline No (Fixed), we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points. This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and a state-of-the-art architecture.
MAEs dominate Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG and S3D) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training. Given the age of Models Genesis - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF. We attribute this to a combination of two factors: 1. Models Genesis was originally published and trained on an outdated 3D-UNet [7] and outside of the powerful nnU-Net framework [20]. This highlights the importance of avoiding Pitfall 2: Training on a state-of-the-art backbone. 2. VoCo and VF were introduced in conjunction with architectures they were optimized for. By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture pre-training combination may be suboptimal for the new CNN backbone.
@@ 5.2 Ablation experiments
To assess the generalization capability of the proposed pre-training method, we tested two scenarios. First, we evaluated fine-tuning our method on an unseen modality using the TOF Angiography Aneurysms dataset (D12). As shown in Tab. 9 without pre-training, the fixed configuration suffers a performance drop of 20 Dice points. We attribute this to the significant difference in median spacing for the downstream task ($[0.50,0.43,0.43]$ mm), which has a higher resolution than the fixed target spacing of $[1,1,1]$ mm used in the pre-training experiments. This lower resolution likely increases the difficulty of segmenting small aneurysms. Despite this decrease, pre-training mitigates some of this degradation and proves highly beneficial compared to training the same configuration from scratch. Interestingly, Models Genesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences. Second, we fine-tuned on the D2 dataset using only the T1 contrast-enhanced (T1ce) sequence and applied these models directly to D13 without any additional fine-tuning. While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that pre-training improves generalization across different centers, with our S3D model yielding the best performance.
@@ 6 Conclusion
7 Acknowledgments
This work was partly funded by the Helmholtz Foundation Model Initiative (HFMI) under the subproject ‘The Human Radiome Project’ (THRP). Part of this work was funded by Helmholtz Imaging (HI), a platform of the Helmholtz Incubator on Information and Data Science.
@@ Appendix B Longer Training schedule
Table 5: Publicly available checkpoint trained on the ABCD dataset: To provide a public available checkpoint, we retrained our proposed model on the ABCD dataset, indicated by a *. It performs slightly worse than the network pre-trained on the private dataset.
@@ C.2 Comparison to previous work on CT data
Table 6: Forty images with SSL are almost as good as all data from-scratch! The pre-trained S3D model almost reaches the performance of the model trained from-scratch with only 40 training cases, with the exception of D4. Overall train/val/test dataset size was 38/10/12 for D1, 67/17/21 for D2, 166/42/52 for D3, 419/105/131 for D4, 134/34/42 for D5. Results in the table are reported on the validation set. full: Uses all train samples of the dataset. * D1 has only 38 training cases for the train split.
@@
Experiment Setting No Dyn. No Fixed VoCo VF MG S3D-B
Modality shift TOF Angio. Aneurysms(D12) 42.61 22.76 22.32 31.21 34.60 28.72
In Distribution Brain Mets (D2) 72.81 67.93 64.34 71.69 69.05 71.56
Patient shift Brain Mets (D13) 64.08 61.61 56.78 63.95 64.22 64.54
Table 10: Fine-tuning length: When initializing from our pre-trained checkpoint, it is possible to achieve a large fraction of the final performance after less than 15% of the normal training time. Despite this a full training schedule reaches better performance. These experiments were conducted using S3D long on the validation splits.
@@
Appendix D Distinction to AMAEs
Concurrently with this work, [25] introduced the AMAEs framework. Like our approach, it utilizes a dataset of approximately 40k 3D images and employs a state-of-the-art CNN architecture. However, in contrast to our more comprehensive evaluation, their assessment is based on three in-distribution and one out-of-distribution downstream dataset.
While evaluation on three in and one out-of-distribution datasets can suffice to draw some insights, their evaluation setup has additional limitations. First, they constrain themselves to a low-data regime and do not assess whether their performance surpasses a default nnU-Net baseline that is trained for 1000 epochs. Second, their fine-tuning strategy is fixed to a single-channel input to align with the pre-trained stem weights. While this would be okay for their method, they apply the same limitation to their baselines, which typically utilize all available modalities, which can lead to stronger performance. For instance, in their evaluation on the BraTS dataset, only one of the four available modalities was used, potentially limiting the effectiveness of the baseline models. These choices may impact result reliability and align with Pitfall 3—unreliable evaluation practices. To address such concerns, our work adopts a very comprehensive evaluation strategy to allow drawing reliable conclusions on the state-of-the-art in self-supervised learning for 3D medical image segmentation for the first time.
Appendix E Acknowledgement of ABCD
Data used in the preparation of this article were obtained from the Adolescent Brain Cognitive DevelopmentSM (ABCD) Study (https://abcdstudy.org), held in the NIMH Data Archive (NDA). This is a multisite, longitudinal study designed to recruit more than 10,000 children age 9-10 and follow them over 10 years into early adulthood. The ABCD Study® is supported by the National Institutes of Health and additional federal partners under award numbers U01DA041048, U01DA050989, U01DA051016, U01DA041022, U01DA051018, U01DA051037, U01DA050987, U01DA041174, U01DA041106, U01DA041117, U01DA041028, U01DA041134, U01DA050988, U01DA051039, U01DA041156, U01DA041025, U01DA041120, U01DA051038, U01DA041148, U01DA041093, U01DA041089, U24DA041123, U24DA041147. A full list of supporters is available at https://abcdstudy.org/federal-partners.html. A listing of participating sites and a complete listing of the study investigators can be found at https://abcdstudy.org/consortium_members/. ABCD consortium investigators designed and implemented the study and/or provided data but did not necessarily participate in the analysis or writing of this report. This manuscript reflects the views of the authors and may not reflect the opinions or views of the NIH or ABCD consortium investigators.